9. A fall in the price of X from Rs. 12 to Rs. 8 causes an increase in the

quantity of Y demanded from 900 to 1,100 units. X and Y are which type of

goods in the question.


1
Expert's answer
2021-02-28T11:38:22-0500

Let's first find the cross elasticity of demand:


"E_{xy}=\\dfrac{\\%\\Delta Q_x}{\\%\\Delta P_y}=\\dfrac{\\dfrac{Q_2-Q_1}{\\dfrac{Q_2+Q_1}{2}}}{\\dfrac{P_2-P_1}{\\dfrac{P_2+P_1}{2}}},\n\n\u200b\t\n \n\u200b""E_{xy}=\dfrac{\dfrac{1100-900}{\dfrac{1100+900}{2}}}{\dfrac{8-12}{\dfrac{8+12}{2}}}=-0.5."

Since "E_{xy}\lt 0" the goods X and Y are complements (complementary goods).